# Non-classical correlations between a quantum probe and complex quantum systems in presence of noise

### v1

@@ 3 Qubit-tardigrade system
The time evolution plot of the negativity measure of each of the light, qubit, and tardigrade subsystems is shown in Fig. 4. Here, the coupling strengths are fixed at $g_{ql}=0.15\times 10^{9}$z and $f_{tl}=0.05\times 10^{9}$z. The noise level is set by fixing the decay rates at $\kappa_{l}=3.7\times 10^{7}$, $\delta_{q}=2.5\times 10^{7}$ and $\gamma_{t}=1.8\times 10^{7}$, respectively. The system is evolved till time, $200ns$, in order to get steady state values for negativity. In this case, each run of code for a fixed set of parameters on a computational node with 28 cores took around 8 hours on average; hence, unlike the case for the previous section, it was more computationally feasible here to do further analysis based on varying coupling strengths and levels of noise, due to the much smaller dimensional structure arising from this model having only one light mode under consideration.
The next analysis is of the steady state values for the negativity of each subsystem, varied for discrete levels of noise and coupling strengths. We vary the decay rates in terms of different orders of magnitude below the respective frequencies of each subsystem. Coupling constants are varied in equal intervals within the range of $[0,0.3]\times 10^{9}$.
Fig. 4(a) plots the time evolution of the purity of the density state for different noise levels, and Fig. 4(b) is a scatter plot of the steady state values of the negativity measure of each subsystem varied against different noise levels. The coupling strengths for both figures are fixed as before. It can be seen that, for noise levels up to two orders of magnitude lower than the frequency order, the purity of the system tends to approach 0.94 asymptotically. However, when the decay rates are closer to the order of magnitude of the frequencies, the system loses little of its purity. Similarly, the steady state values of the negativity of each subsystem appear to be constant until the decay rates get closer to the order of the frequencies, when they sharply decrease.
Finally, in Fig. 6, three sets of scatter plots show the behavior of the steady state values of negativity for each subsystem with respect to the coupling strengths for different noise levels. In each row, the light-qubit coupling, $g_{ql}$, is varied in the first scatter plot, keeping $f_{tl}$ fixed. In the second plot on each row, the light-tardigrade coupling, $f_{tl}$, is varied, and finally, in the third scatter plot, both couplings are varied while keeping them equal to each other. We observe that the steady state values of the detectable entanglement (particularly that between the light modes) increase with increase in the relevant couplings while they go down with increasing noise as expected.
@@ 4 Conclusions
We have examined in detail the interesting idea put forward in Krisnanda2018 that non-classical correlations induced in a microscopic quantum probe by its interaction with a macroscopic system can be used to detect the presence of quantum features in the macroscopic system. We focused on a case where the quantum probe is continually interacting with the macroscopic system, and at the same time, both are subject to environmental noise and decoherence. Our objective was to explore the robustness of this way of detecting quantum features in the macroscopic system against noise, which is to be expected in most realistic scenarios. In place of the heuristic models for noise used in previous works along similar lines, we have used a GKSL master equation to describe the open dynamics of the probe and macroscopic system. We numerically integrated the relevant master equations using realistic parameter values and observed that, indeed, entanglement and other non-classical correlations are built up between the quantum probe and various components of the macroscopic system. We also verified that these correlations are robust against noise and saturate to finite values for moderate and low levels of noise in the system.

### v2

@@ 3 Qubit-tardigrade system
The time evolution plot of the negativity measure of each of the light, qubit, and Tardigrade subsystems in the presence of decay and dephasing is shown in Fig. 4. Here, the coupling strengths are fixed at $g_{ql}=0.15\times 10^{9}$ Hz and $f_{tl}=0.05\times 10^{9}$ Hz. The noise level is set by fixing the decay rates at $\kappa_{l}=\tilde{\kappa}_{l}=3.7\times 10^{7}$ Hz, $\delta_{q}=\tilde{\delta}_{q}=2.5\times 10^{7}$ Hz, and $\gamma_{t}=\tilde{\gamma}_{t}=1.8\times 10^{7}$ Hz, respectively. The system is evolved till time, $200ns$, in order to get steady state values for negativity. In this case, each run of code for a fixed set of parameters on a computational node with 28 cores took around 6 hours on average; hence, unlike the case for the previous section, it was more computationally feasible here to do further analysis based on varying coupling strengths and levels of noise, due to the much smaller dimensional structure arising from this model having only one light mode under consideration.
The next analysis is of the steady state values for the negativity of each subsystem, varied for discrete levels of noise (both decay and dephasing) and coupling strengths. We vary the decay rates in terms of different orders of magnitude below the respective frequencies of each subsystem. Coupling constants are varied in equal intervals within the range of $[0,0.3]\times 10^{9}$ Hz.
Fig. 5(a) plots the time evolution of the purity of the density state for different noise levels, and Fig. 5(b) is a scatter plot of the steady state values of the negativity measure of each subsystem varied against different noise levels. The coupling strengths for both figures are fixed as before. It can be seen that, for noise levels up to two orders of magnitude lower than the frequency order, the purity of the system tends to approach 0.78 asymptotically. However, when the decay rates are closer to the order of magnitude of the frequencies, the system loses little of its purity. Similarly, the steady state values of the negativity of each subsystem appear to be constant until the decay rates get closer to the order of the frequencies, when they sharply decrease.
Finally, in Fig. 6, three sets of scatter plots show the behavior of the steady state values of negativity for each subsystem with respect to the coupling strengths for different levels of noise (both decay and dephasing). In each row, the light-qubit coupling, $g_{ql}$, is varied in the first scatter plot, keeping $f_{tl}$ fixed. In the second plot on each row, the light-tardigrade coupling, $f_{tl}$, is varied, and finally, in the third scatter plot, both couplings are varied while keeping them equal to each other. We observe that the steady state values of the detectable entanglement (particularly that between the light modes) increase with increase in the relevant couplings while they go down with increasing noise as expected.
@@ 4 Conclusions
We have examined in detail the interesting idea put forward in [15] that non-classical correlations induced in a microscopic quantum probe by its interaction with a macroscopic system can be used to detect the presence of quantum features in the macroscopic system. We focused on a case where the quantum probe is continually interacting with the macroscopic system, and at the same time, both are subject to environmental noise and decoherence. Our objective was to explore the robustness of this way of detecting quantum features in the macroscopic system against noise, which is to be expected in most realistic scenarios. In place of the heuristic models for noise used in previous works along similar lines, we have used a GKSL master equation to describe the open dynamics of the probe and macroscopic system. We numerically integrated the relevant master equations using realistic parameter values and observed that, indeed, entanglement and other non-classical correlations are built up between the quantum probe and various components of the macroscopic system. We also verified that these correlations are robust against noise and saturate to finite values for moderate and low levels of noise in the system.